Need Help badly. Asked for help and didn't get any.
3) For the equation x - 2√x = 0, perform the following:
a) Solve for all values of x that satisfies the equation.
Answer:
Show work in this space.




b) Graph the functions y = x and y = 2√x on the same graph ( by plotting points if necessary). Show the points of intersection of these two graphs.
Graph:
Points of Intersection:

Answer by fractalier(6550)

For x - 2√x = 0, we move the root to the other side and then square both sides:
x = 2√x
x^2 = 4x
x^2 - 4x = 0
x(x - 4) = 0 and
x = 4 or x = 0
If we were to graph the two functions, we would find that they intersect at the points (0, 0) and (4, 4).